# RL in the Wild: Characterizing RLVR Training in LLM deployment

### v1

@@ 1 Introduction
Within this broader context of LLM development, RLVR has emerged as a critical methodology for post-training optimization. Unlike conventional pretraining[9] and SFT[64, 4, 47] phases, RLVR introduces a fundamentally different computational paradigm where models interact with environments, receive feedback signals, and iteratively refine their behavior through policy optimization algorithms such as Proximal Policy Optimization (PPO) [57] and Group Relative Policy Optimization (GRPO) [58]. This approach has proven essential for enhancing their reasoning capabilities[62], leading to widespread adoption in both research and industry.
@@
(1) Longer step span. In RLVR, a single training step often takes on the order of minutes and, in some cases, can exceed one hour. By contrast, other LLM training tasks typically complete a step in several seconds, while an inference step (prefill and decode) is usually under one second. This significantly increases fault recovery costs, hinders real-time monitoring.
(2) Complicated workflow. In RLVR, a single training run requires multiple models to collaborate according to the algorithmic logic, and is often divided into several stages, such as rollout, inference, and training. By contrast, other LLM training tasks typically involve only one model and a single training stage[47, 53]. This introduces challenges to resource allocation and handling failure propagation across stages.
(3) Unknown sequence length. In prior LLM training tasks, the workload was known in advance, enabling offline reordering and scheduling policy selection[11]. In RLVR, however, the workload—much like that of LLM serving—cannot be predetermined and exhibits dynamics and abrupt shifts. This leads to suboptimal parallelization strategies and difficulty in applying appropriate workload-aware scheduling.
@@
To bridge this gap, we present PolyTrace and an in-depth analysis of RLVR training systems based on our operational experiences across multiple production tasks. To the best of our knowledge, this represents the first comprehensive study of RLVR workloads in real-world deployment scenarios. Our analysis draws upon extensive traces collected from large-scale RLVR training jobs, encompassing detailed profiling data, resource utilization metrics, and system performance logs. Our further key findings can be summarized as follows:
@@
We believe the observations and insights from our work are broadly applicable. Based on our analysis, we highlight the profound impact of workload distribution on the performance of RL training systems. Consequently, our traces and benchmark suite are publicly available at https://github.com/Zhou-jiecheng/PolyTrace to facilitate further research and development in this area.
@@ 2.1 LLM Reinforcement Learning
1. Rollout. The actor model, analogous to an agent model in the RL setting, performs autoregressive generation[70] starting from an initial state. Each generated token constitutes an action that transitions to a new state. Once generation terminates, the resulting sequence is treated as a trajectory for subsequent stages. This stage is usually executed by inference engines (e.g., vLLM[28], SGLang[86], and LMdeploy[82]). The actor model may also invoke external tools (e.g., image processing[88], web search[24]) at this stage.
@@
Because training objectives differ across tasks [62], models "react" differently to each task, resulting in highly diverse workloads. In such a multi-stage pipeline, workload characteristics exert a complex and pronounced impact on the system.
@@ 3.1 Vision-Task Workloads
Figure 7: Similarity of distribution across taining steps of Image Understanding Task. The metric in the figure is 1 minus the Jensen–Shannon (JS) divergence.

### v2

@@ 1 Introduction
Within this broader context of LLM development, RLVR has emerged as a critical methodology for post-training optimization. Unlike conventional pretraining[9] and SFT[64, 4, 47], RLVR introduces a fundamentally different computational paradigm where models interact with environments, receive rewards, and iteratively refine their behavior through policy optimization algorithms such as Proximal Policy Optimization (PPO) [57] and Group Relative Policy Optimization (GRPO) [58]. This approach has proven essential for enhancing their reasoning capabilities[62], leading to widespread adoption in both research and industry.
@@
(1) Longer step span. In RLVR, a single training step often takes on the order of minutes and, in some cases, can exceed one hour. By contrast, other LLM training tasks typically complete a step in several seconds, while an inference step (prefill and decode) is usually under 1s.
(2) Complicated workflow. In RLVR, a single training run requires multiple models to collaborate according to the algorithmic logic, and is often divided into several stages, such as rollout, inference, and training. By contrast, other LLM training tasks typically involve only one model and a single training stage[47, 53]. This introduces challenges to resource allocation and handling error propagation.
(3) Unknown sequence length. In prior LLM training tasks, the workload was known in advance, enabling offline reordering and scheduling policy selection[11]. In RLVR, however, the workload—much like that of LLM serving—cannot be predetermined and exhibits dynamics and abrupt shifts. This leads to suboptimal parallelization and difficulty in applying appropriate scheduling strategies.
@@
To bridge this gap, we present PolyTrace and an in-depth analysis of RLVR training systems based on our operational experiences across multiple production tasks. To the best of our knowledge, this represents the first comprehensive study of RLVR workloads in real-world deployment scenarios. Our analysis draws upon extensive traces collected from large-scale RLVR training jobs, encompassing detailed profiling data and system performance logs. Our further key findings can be summarized as follows:
@@
[…]
@@ 2.1 LLM Reinforcement Learning
1. Rollout. The actor model, analogous to an agent model in the RL setting, performs autoregressive generation[70] starting from an initial state. Each generated token constitutes an action that transitions to a new state. Once generation terminates, the resulting sequence is treated as a trajectory for subsequent stages. This stage is usually executed by inference engines (e.g., vLLM[28], SGLang[86], and LMdeploy[82]). The actor model may also invoke external tools (e.g., image processing[88]) in rollout.
@@
Because training objectives differ across tasks [62], models "react" differently to each task, resulting in highly diverse workloads. In such a multi-stage pipeline, workload characteristics exert a pronounced impact on the system.
@@ 3.1 Vision-Task Workloads
Figure 7: Similarity of distribution across training steps of Image Understanding Task. The metric in the figure is 1 minus the Jensen–Shannon (JS) divergence.